# Arithmetic Control of LLMs for Diverse User Preferences: Directional Preference Alignment with Multi-Objective Rewards

### 29 Feb 2024 v2

@@ Abstract
Fine-grained control over large language models (LLMs) remains a significant challenge, hindering their adaptability to diverse user needs. While Reinforcement Learning from Human Feedback (RLHF) shows promise in aligning LLMs, its reliance on scalar rewards often limits its ability to capture diverse user preferences in real-world applications. To address this limitation, we introduce the Directional Preference Alignment (DPA) framework. Unlike the scalar-reward RLHF, DPA incorporates multi-objective reward modeling to represent diverse preference profiles. Additionally, DPA models user preferences as directions (i.e., unit vectors) in the reward space to achieve user-dependent preference control. Our method involves training a multi-objective reward model and then fine-tuning the LLM with a preference-conditioned variant of Rejection Sampling Finetuning (RSF), an RLHF method adopted by Llama 2. This method enjoys a better performance trade-off across various reward objectives. In comparison with the scalar-reward RLHF, DPA offers users intuitive control over LLM generation: they can arithmetically specify their desired trade-offs (e.g., more helpfulness with less verbosity). We also validate the effectiveness of DPA with real-world alignment experiments on Mistral-7B. Our method provides straightforward arithmetic control over the trade-off between helpfulness and verbosity while maintaining competitive performance with strong baselines such as Direct Preference Optimization (DPO). The code and trained model are released at github.com/Haoxiang-Wang/directional-preference-alignment.
@@ 1 Introduction
While recent advancements in RLHF are noteworthy, a fundamental challenge persists due to problem misspecification. This means that a single reward function may not sufficiently capture complex human values. For example, a generative model aligned by RLHF for helpfulness tends to produce verbose responses as shown in Figure 1 (Left) Singhal et al. (2023b), even though many users prefer answers that are both helpful and concise. Assuming ae-objective reward implies a total order over preferences, which is hard to satisfy when the preference is aggregated across a diverse set of human groups (May, 1954; Tversky, 1969), because humans typically have a set of intricate or even contradictory targets (Biyik and Sadigh, 2018). In real-world applications, the scalar-reward RLHF tends to align the LLMs toward an “average-user” preference, which cannot capture the complicated nature of human preferences and can be unfair for the under-represented groups (Feffer et al., 2023). For example, consider User-1, 2, 3, and responses $A$, $B$, $C$ in Fig. 2 (Left). User-1 and 3 prefer response $B$ over $C$ ($B\prec C$), while User-2 prefers $C$ over $B$ ($C\prec B$). This could occur as response $C$ is more verbose than $B$, while User-2 prefers concise answers. When these diverse preferences are aggregated across human groups, the typical reward models with scalar rewards tend to learn the “average-user” preference (which is $B\prec C$ in this case), overlooking the individual preference of User-2, as shown in Figure 2 (Middle). This is also known as the “Condorcet paradox” in the theory of social choice Gehrlein (2002). In general, human opinions and expertise can vary significantly (Coello, 2000; Bobu et al., 2023; Bansal et al., 2023). Meanwhile, the importance of these targets may also change over time, depending on the users and their expectations.
@@ 2.3 Discussion with Existing Methods
Comparison with SteerLM Dong et al. (2023b). Recall that we have multi-objective reward $r=\left<r_{1},r_{2},...,r_{k}\right>$ of each response $y$ to the prompt $x$. Dong et al. (2023b) first fine-tunes the generative model to maximize the likelihood of $y$ by taking both $x$ and $r$ as the input prompts:
@@
Comparison with Soup Methods Rame et al. (2023); Jang et al. (2023). Soup methods trains a policy $\theta_{i}$ for each reward objective. Let $r_{i}(x,y)$ denote the $i$-th objective, we have:
@@ 4 Related Works
The landscape of natural language processing has been profoundly transformed in recent years through the development of large language models (LLMs), showcasing human-level proficiency across a range of tasks including text classification, generation, and complex reasoning. This progress stems from extensive pre-training on vast datasets, enabling these models to address diverse challenges. Despite their achievements, a distinction arises between closed-source models (e.g., GPT-3 (Brown et al., 2020), Bard (Google, 2023), Claude (Anthropic, 2023), and PaLM (Chowdhery et al., 2023)), often surpassing their open-source counterparts (e.g., megatron-turing-530b (Smith et al., 2022), and Bloom (Workshop et al., 2022)) in performance (Liang et al., 2022), which poses challenges for open-source research. However, initiatives like Meta’s LLaMA (Touvron et al., 2023) and subsequent works such as Alpaca (Taori et al., 2023), Vicuna (Chiang et al., 2023), and LMFlow (Diao et al., 2023), demonstrate significant open-source contributions that continue to push the boundaries of what’s possible with LLMs. These advancements enabled by the fine-tuning techniques, aim to improve LLMs’ ability and adapt to a wide range of domains and tasks. Nonetheless, as these generative foundation models advance, they still face problems like implicit biases, underscoring the need for ongoing alignment and ethical considerations in their development and application. In this paper, we focus on how to align LLMs with human preferences, including the principles of being helpful, honest, and harmless as outlined by (Askell et al., 2021). This procedure is often achieved by Reinforcement Learning with Human Feedback (RLHF) Ouyang et al. (2022).

### 06 Mar 2024 v3

@@ Abstract
Fine-grained control over large language models (LLMs) remains a significant challenge, hindering their adaptability to diverse user needs. While Reinforcement Learning from Human Feedback (RLHF) shows promise in aligning LLMs, its reliance on scalar rewards often limits its ability to capture diverse user preferences in real-world applications. To address this limitation, we introduce the Directional Preference Alignment (DPA) framework. Unlike the scalar-reward RLHF, DPA incorporates multi-objective reward modeling to represent diverse preference profiles. Additionally, DPA models user preferences as directions (i.e., unit vectors) in the reward space to achieve user-dependent preference control. Our method involves training a multi-objective reward model and then fine-tuning the LLM with a preference-conditioned variant of Rejection Sampling Finetuning (RSF), an RLHF method adopted by Llama 2. This method enjoys a better performance trade-off across various reward objectives. In comparison with the scalar-reward RLHF, DPA offers users intuitive control over LLM generation: they can arithmetically specify their desired trade-offs (e.g., more helpfulness with less verbosity). We also validate the effectiveness of DPA with real-world alignment experiments on Mistral-7B. Our method provides straightforward arithmetic control over the trade-off between helpfulness and verbosity while maintaining competitive performance with strong baselines such as Direct Preference Optimization (DPO). The code and trained model are released at https://github.com/Haoxiang-Wang/directional-preference-alignment.
@@ 1 Introduction
While recent advancements in RLHF are noteworthy, a fundamental challenge persists due to problem misspecification. This means that a single reward function may not sufficiently capture complex human values. For example, a generative model aligned by RLHF for helpfulness tends to produce verbose responses as shown in Figure 1 (Left) (Singhal et al., 2023b), even though many users prefer answers that are both helpful and concise. Assuming scalar-objective reward implies a total order over preferences, which is hard to satisfy when the preference is aggregated across a diverse set of human groups (May, 1954; Tversky, 1969), because humans typically have a set of intricate or even contradictory targets (Biyik and Sadigh, 2018). In real-world applications, the scalar-reward RLHF tends to align the LLMs toward an “average-user” preference, which cannot capture the complicated nature of human preferences and can be unfair for the under-represented groups (Feffer et al., 2023). For example, consider User-1, 2, 3, and responses $A$, $B$, $C$ in Fig. 2 (Left). User-1 and 3 prefer response $B$ over $C$ ($B\prec C$), while User-2 prefers $C$ over $B$ ($C\prec B$). This could occur as response $C$ is more verbose than $B$, while User-2 prefers concise answers. When these diverse preferences are aggregated across human groups, the typical reward models with scalar rewards tend to learn the “average-user” preference (which is $B\prec C$ in this case), overlooking the individual preference of User-2, as shown in Figure 2 (Middle). This is also known as the “Condorcet paradox” in the theory of social choice (Gehrlein, 2002). In general, human opinions and expertise can vary significantly (Coello, 2000; Bobu et al., 2023; Bansal et al., 2023). Meanwhile, the importance of these targets may also change over time, depending on the users and their expectations.
@@ 2.3 Discussion with Existing Methods
Comparison with SteerLM (Dong et al., 2023b). Recall that we have multi-objective reward $r=\left<r_{1},r_{2},...,r_{k}\right>$ of each response $y$ to the prompt $x$. Dong et al. (2023b) first fine-tunes the generative model to maximize the likelihood of $y$ by taking both $x$ and $r$ as the input prompts:
@@
Comparison with Soup Methods (Rame et al., 2023; Jang et al., 2023). Soup methods trains a policy $\theta_{i}$ for each reward objective. Let $r_{i}(x,y)$ denote the $i$-th objective, we have:
@@ 4 Related Works
The landscape of natural language processing has been profoundly transformed in recent years through the development of large language models (LLMs), showcasing human-level proficiency across a range of tasks including text classification, generation, and complex reasoning. This progress stems from extensive pre-training on vast datasets, enabling these models to address diverse challenges. Despite their achievements, a distinction arises between closed-source models (e.g., GPT-3 (Brown et al., 2020), Bard (Google, 2023), Claude (Anthropic, 2023), and PaLM (Chowdhery et al., 2023)), often surpassing their open-source counterparts (e.g., megatron-turing-530b (Smith et al., 2022), and Bloom (Workshop et al., 2022)) in performance (Liang et al., 2022), which poses challenges for open-source research. However, initiatives like Meta’s LLaMA (Touvron et al., 2023) and subsequent works such as Alpaca (Taori et al., 2023), Vicuna (Chiang et al., 2023), and LMFlow (Diao et al., 2023), demonstrate significant open-source contributions that continue to push the boundaries of what’s possible with LLMs. These advancements enabled by the fine-tuning techniques, aim to improve LLMs’ ability and adapt to a wide range of domains and tasks. Nonetheless, as these generative foundation models advance, they still face problems like implicit biases, underscoring the need for ongoing alignment and ethical considerations in their development and application. In this paper, we focus on how to align LLMs with human preferences, including the principles of being helpful, honest, and harmless as outlined by (Askell et al., 2021). This procedure is often achieved by Reinforcement Learning with Human Feedback (RLHF) (Ouyang et al., 2022).